# CTP-SCU/2025011

### v1

@@ Abstract
We provide a field-theoretic method to calculate entanglement entropy of CFT in all dimensions, for arbitrary shape of entangling surfaces. Ryu-Takayanagi formula is manifested in this formalism.
@@ I Introduction
Entanglement properties of quantum systems are playing an increasingly crucial role in physics. An intrinsic quantity counting the degrees of freedom of entanglement is the von Neumann entropy, also known as entanglement entropy. It is defined by $S_{A}=-\mathrm{Tr}\,\rho_{A}\log\rho_{A}$ for a reduced density matrix $\rho_{A}$ associated with a subsystem $A$, where the total system is bipartitioned into two subsystems, $A$ and its complement $A^{c}$. In quantum field theory (QFT), one usually considers the entropy of entanglement between two adjacent and complementary subregions $A$ and $A^{c}$, denoted as $S(A:A^{c})$. Although the calculation of entanglement entropy in two dimensional CFT has been well established using the replica trick [1, 2], no approach currently exists for performing such calculations in higher dimensional CFTD for $D>2$.
The entropy $S(A:A^{c})$ typically suffers an ultraviolet (UV) divergence due to the very intense entanglement between contiguous fields. This divergent behavior renders it extremely challenging, if not entirely impossible, to derive exact relations between QFT and gravity. A natural question arises: are there any finite elementary entanglement entropies? For such quantities to exist, two conditions should be satisfied:
@@
It is not hard to understand that the annular CFT2 in Figure 1 precisely meets these two conditions. In prior work with collaborators [3, 4, 5], we demonstrated that the entanglement entropy $S(A:B)$ between disjoint intervals $A$ and $B$ is indeed finite, for static, covariant and thermal configurations, respectively. Notably, $S(A:B)$ can be equivalently interpreted as the entanglement entropy of the mixed state $\rho_{AB}$ in the infinite system, as illustrated in Figure 2. The well-known divergent entanglement entropies for adjacent intervals emerge naturally as the adjacent limits of disjoint $S(A:B)$.
Figure 1: The Euclidean pure state density matrix for the annular CFT2. Two distinct intervals $A$ and $B$ have finite entanglement entropy $S(A:B)$.
Figure 2: After subtracting segments $C$ and $D$ with two discs in the infinite system, we obtain an annular region in which $A$ and $B$ are in a pure entangled state $\psi_{AB}$. It is identical to the annular CFT2.
A critical insight is the straightforward generalization of the annular CFT2 to higher dimensions: CFTD on a solid torus $\mathbb{B}^{D-1}\times S^{1}$, see Figure 3, where $\mathbb{B}^{D-1}$ is a $(D-1)$-dimensional ball. Since the replica of a solid torus is still a solid torus, calculating entanglement entropies simplifies to computing the partition function on this solid torus! The usual divergent entanglement entropies for adjacent regions are again easily obtained by taking adjacent limits. In this paper, we only study the ball shaped entangling region. But one can easily see that this framework works for entangling surfaces of arbitrary shape.
For $D>2$, the relationship between the solid torus CFTD and the infinite CFTD on $\mathbb{R}^{D}$ mirrors the $D=2$ case. Consider calculating the entanglement entropy between two codimension-one disjoint discs $A$ and $B$ in $\mathbb{R}^{D}$, corresponding to the mixed state $\rho_{AB}$ of the infinite system. Analogous to the two-dimensional scenario, removing a D-dimensional solid torus between $A$ and $B$ yields also a D-dimensional solid torus depicted in Figure 3 — precisely the SUBTRACTION procedure introduced in [3].
The remainder of this paper is outlined as follows. In section II, we derive the explicit expression of the entanglement entropy between two disjoint balls in all dimensions. In section III, by taking the adjacent limits, the divergent entanglement entropies for adjacent balls in all dimensions are obtained, where the area law [6, 7] is manifested. In section IV, we provide examples in CFT2 and CFT4. In section V, we show how RT formula [8, 9] naturally emerges in this framework and discuss arbitrary shape of entangling surfaces. In section VI, we find perfect agreement of our results with the holographic calculation.
@@ II Entanglement entropy between two disjoint balls
For a CFT in $D$-dimensional Euclidean spacetime with the metric $\mathrm{d}s_{0}^{2}=\mathrm{d}t_{\text{E}}^{2}+\mathrm{d}y^{2}+\underset{i=1}{\overset{D-2}{\sum}}\mathrm{d}x_{i}^{2}$, we consider quantum fields living in a bounded region,
:=$\mathcal{B}:=\left\{(t_{\text{E}},y,x)|\left(\sqrt{t_{\text{E}}^{2}+y^{2}}-\frac{l_{2}+l_{1}}{2}\right)^{2}+\sum_{i=1}^{D-2}x_{i}^{2}<\left(\frac{l_{2}-l_{1}}{2}\right)^{2}\right\},$ (1)
with $l_{2}>l_{1}>0$. At the time slice $t_{\text{E}}=0$, $\mathcal{B}$ consists of two disjoint $(D-1)$-balls $A$ and $B$ with the same radius, as shown in Figure 3. The bounded region $\mathcal{B}$ is intuitively a solid torus $\mathbb{B}^{D-1}\times S^{1}$, where $\mathbb{B}^{D-1}$ is a $(D-1)$-dimensional ball. Up to a Weyl transformation $\mathrm{d}s_{0}^{2}\rightarrow\mathrm{d}s_{0}^{2}/r^{2}$, the spacetime metric becomes
$\mathrm{d}s^{2}=\mathrm{d}\theta^{2}+\left(\mathrm{d}r^{2}+\sum_{i=1}^{D-2}\mathrm{d}x_{i}^{2}\right)/r^{2},$ (2)
in terms of coordinates transformations, $t_{\text{E}}=r\sin\theta$, $y=r\cos\theta$. In this parameterization, we can treat $\theta$ as imaginary time with a periodicity $2\pi$, and the CFT living in $\mathcal{B}$ can be thought as a thermal field theory with the inverse temperature $\beta=2\pi$. The partition function of the CFT can be formally written as a trace over the Hilbert space:
$Z=\mathrm{Tr}\,e^{-2\pi H}.$ (3)
The Hamiltonian operator $H$ conjugate to the imaginary time $\theta$ is defined by
@@
where $T$ is the stress-energy tensor.
Figure 3: Entangling regions $A$ and $B$ (colored regions) in the bounded region $\mathcal{B}$ (blue region).
Starting point of our computations is the observation that the vacuum entanglement entropy between $A$ and $B$ can be computed by the replica trick [1]. Such techniques allow us to write the Rényi entropy as
$S^{(n)}=\frac{1}{1-n}\log\left[\frac{Z_{\mathcal{B}_{n}}}{Z^{n}}\right],$ (4)
where $\mathcal{B}_{n}$ is made by cyclically gluing $n$ copies of $\mathcal{B}$ with a cut along $A$ (or equivalently $B$), and the metric of $\mathcal{B}_{n}$ might be given by
$\mathrm{d}s_{(n)}^{2}=n^{2}\mathrm{d}\theta^{2}+\left(\mathrm{d}r^{2}+\sum_{i=1}^{D-2}\mathrm{d}x_{i}^{2}\right)/r^{2}.$ (5)
The partition function of CFT living on $\mathcal{B}_{n}$ is then
@@
Then, the Rényi entropy becomes
$S^{(n)}=\frac{1}{1-n}\log\left[\frac{e^{-2\pi\langle H\rangle/n}}{e^{-2n\pi\langle H\rangle}}\right]=-2\pi\left(1+\frac{1}{n}\right)\langle H\rangle.$ (9)
The entropy $S(A:B)$ of entanglement between two disjoint balls $A,B$ therefore reads
$S(A:B)=S^{(1)}=-4\pi\langle H\rangle.$ (10)
@@
The entanglement entropy $S(A:B)$ simply depends on the ground state energy $\langle H\rangle$ localized in a spatial ball $\mathbb{B}^{D-1}$:
@@
where $\mathcal{E}_{\text{vac}}$ is the Casimir energy density that is usually a negative constant in our discussions. Consequently, the entanglement entropy $S(A:B)$ is given by
$S(A:B)$ $=$ $-4\pi\mathcal{E}_{\text{vac}}\mathrm{Vol}(\mathbb{B}^{D-1}),$ (12)
@@
with the standard hypergeometric function $\,{}_{2}F_{1}$. The parameters $l_{1}$ and $l_{2}$ are the inner and outer radii of the solid torus, as indicated in the left panel of Figure 5.
Figure 4: Entangling regions (colored regions) are bounded by ($D-2$)-spheres (black circles), respectively. Left panel (the cavity configuration): A ($D-1$)-ball living in a spheric cavity is entangled with the region outside the cavity. Right panel (the juxtaposed configuration): Two disjoint ($D-1$)-balls with different radii are entangled.
Generally, there exist two different configurations of our interest, namely, the cavity configuration and the juxtaposed configuration illustrated in Figure 4. Each configuration is uniquely fixed by boundaries of $A,B$, i.e., two disjoint ($D-2$)-spheres. Every ($D-1$)-ball is bounded by a ($D-2$)-sphere $\Sigma(\vec{x},r)$ located at $\vec{x}\in\mathbb{R}^{D-1}$ with the radius $r$, respectively. An elegant substitute for conformality is the inversive product $\varrho$ of two spheres $\Sigma(\vec{x},r)$ and $\Sigma(\vec{x}^{\prime},r^{\prime})$ [10]
@@
In the case that two disjoint ($D-1$)-balls $A,B$ share the same radius, we have
@@
The inversive product $\varrho$ is a conformal invariant, as it remains invariant under global conformal transformations. If two configurations have different inversive products, they cannot be conformally mapped onto each other. The set of all configurations falls into classes of conformally equivalent regions, with each class characterized by the inversive product $\varrho$ of that class. Consequently, any two configurations sharing the same $\varrho$ must be conformally equivalent. To compute the entanglement entropy $S(A:B)$ for any two disjoint spheres in general configurations, it suffices to evaluate the inversive product $\varrho$ of the spheres. Thus, we conclude that the entanglement entropy between any two disjoint spheres is given by [^2] For the time dependent covariant scenario, simply replace $\vec{x}$ with $x^{\mu}$.
$S=-4\mathcal{E}_{\text{vac}}\pi^{\frac{D}{2}}\frac{\Gamma\left(\frac{D}{2}\right)}{\Gamma\left(D\right)}\left(\frac{2\sqrt{2}}{\sqrt{\varrho+1}-\sqrt{2}}\right)^{D-1}\,_{2}F_{1}\left(D-1,\frac{D}{2};D;\frac{2\sqrt{2}}{\sqrt{2}-\sqrt{\varrho+1}}\right).$ (15)
@@ III Emergent area law in various dimensions
Figure 5: Left panel: The time zero slice of $\mathcal{B}$ in $D$ dimensions, where $A$ and $B$ are two disjoint ($D-1$)-balls with the same radius (colored regions) living at $t_{\text{E}}=0$. Right panel: The ($D-1$)-ball $A$ with the radius $R-\epsilon$ is in a spheric cavity with the radius $R+2\epsilon$, and the region $B$ is outside the cavity. Here, $\epsilon$ plays a role of the UV cutoff.
We first consider the entanglement entropy $S(A:B)$ for two disjoint ($D-1$)-balls $A,B$ with the same radius in the juxtaposed configuration, as shown in Figure 5. It would be convenient to write down $S(A:B)$ in various dimensions:
@@
Then, we return to the general $D$-dimensional cavity configuration depicted in the right panel of Figure 5, where $A$ is almost adjacent to $B$, more precisely implying that
@@
Plugging $\varrho$ into Equation (16), we obtain the corresponding entanglement entropy in various dimensions:
@@
$D=6:$ $\frac{\pi^{3}\mathcal{E}_{6}}{3}\left(\frac{8R^{4}}{\epsilon^{4}}-\frac{24R^{2}}{\epsilon^{2}}+9+12\log\frac{2R}{\epsilon}\right)+\mathcal{O}\left(\epsilon\right).$ (19)
In $D=2$, the result is the well-known $\log$-law for entanglement entropy. In $D\geq 3$, we can immediately observe the area law of entanglement entropy, which arises as a special limit of our general consideration. We emphasize that in even dimensions, the finite term is physically irrelevant since they can be absorbed into the UV cutoff $\epsilon$. But in odd dimensions, finite terms of entanglement entropies become scheme-independent and play a crucial role in the F-theorem [11, 12, 13]. This is manifested by the exactness of our expressions for entanglement entropy in odd dimensions.

### v2

@@ Abstract
We provide a field-theoretic method to calculate entanglement entropy of CFT in all dimensions. This method works for entangling surfaces of arbitrary shape. The formalism manifests a field-theoretic proof of the Ryu-Takayanagi formula.
@@ I Introduction
Entanglement properties of quantum systems are playing an increasingly crucial role in physics. An intrinsic quantity counting the degrees of freedom of entanglement is the von Neumann entropy, also known as entanglement entropy. It is defined by $S_{A}=-\mathrm{Tr}\,\rho_{A}\log\rho_{A}$ for a reduced density matrix $\rho_{A}$ associated with a subsystem $A$, where the total system is bipartitioned into two subsystems, $A$ and its complement $A^{c}$. In QFT, one usually considers the entanglement entropy between two adjacent and complementary subregions $A$ and $A^{c}$, denoted as $S_{\text{adj}}(A:A^{c})$, where the subscript “adj” stands for adjacent entangling regions. Although the calculation of entanglement entropy in two dimensional CFT has been well established using the replica trick [1, 2], no general approach currently exists for performing such calculations in higher dimensional CFTD for $D>2$.
The entropy $S_{\text{adj}}(A:A^{c})$ typically suffers an ultraviolet (UV) divergence due to the very intense entanglement between contiguous fields. This divergent behavior renders it extremely challenging, if not entirely impossible, to derive exact relations between QFT and gravity. However, for a UV-complete theory such as a CFT, all physical quantities must possess well-defined UV-finite forms. Divergences and regulators, rather than being intrinsic components of the theory itself, should only emerge as artifacts of specific limiting procedures. A natural question thus arises: Do there exist any finite elementary entanglement entropies in CFT? Here, the term elementary is used in contrast to compound quantities such as mutual information, relative entropy, and the like — which are constructed from combinations of elementary entanglement measures. For such finite elementary entanglement entropies to exist, two conditions should be satisfied:
@@
The annular CFT2 in Figure 1 precisely meets these two conditions. Note the time direction is upward. The lower half annulus corresponds to $|\psi\rangle$, and the upper half annulus corresponds to $\langle\psi|$. The density matrix of this pure state is $\rho=|\psi\rangle\langle\psi|$. In prior collaborations [3, 4, 5], we demonstrated that the entanglement entropy $S_{\text{disj}}(A:B)$ between disjoint intervals $A$ and $B$ in this annular CFT is indeed finite, for static, covariant and thermal configurations, respectively. The subscript “disj” denotes disjoint entangling regions. Notably, referring to Figure 2, as shown in [3, 4], $S_{\text{disj}}(A:B)$ can be equivalently interpreted as the entanglement entropy between two disjoint intervals in the infinite system. The well-known divergent entanglement entropies $S_{\text{adj}}(A:B)$ for adjacent intervals emerge naturally as the adjacent limits of disjoint $S_{\text{disj}}(A:B)$.
Figure 1: The Euclidean pure state density matrix $\rho=|\psi\rangle\langle\psi|$ for the annular CFT2. Two distinct intervals $A$ and $B$ have finite entanglement entropy $S(A:B)$.
Figure 2: As shown in [3, 4], after subtracting segments $C$ and $D$ with two discs in the infinite system, we obtain an annular region in which AA and $B$ are in a pure entangled state $\psi_{AB}$. It is identical to the annular CFT2 in Figure 1.
A critical insight is the straightforward generalization of the annular CFT2 to higher dimensions: CFTD on a solid torus $\mathbb{B}^{D-1}\times S^{1}$, see Figure 3, where $\mathbb{B}^{D-1}$ is a (D1)(D-1)-dimensional ball. Since the replica of a solid torus is still a solid torus, calculating entanglement entropies simplifies to computing the partition function on this solid torus! The usual divergent entanglement entropies for adjacent regions are again easily obtained by taking adjacent limits. In this paper, we only study the ball shaped entangling region. But one can easily see that this framework works for entangling surfaces of arbitrary shape.
For $D>2$, the relationship between the solid torus CFTD and the infinite CFTD on $\mathbb{R}^{D}$ mirrors the $D=2$ case. Consider calculating the entanglement entropy between two codimension-one disjoint discs $A$ and $B$ in $\mathbb{R}^{D}$. Analogous to the two-dimensional scenario, removing a $D$-dimensional solid torus between $A$ and $B$ yields also a $D$-dimensional solid torus depicted in Figure 3 — precisely the SUBTRACTION procedure introduced in [3].
The remainder of this paper is outlined as follows. In section II, we derive the explicit expression of the entanglement entropy between two disjoint balls in all dimensions. In section III, by taking the adjacent limits, the divergent entanglement entropies for adjacent balls in all dimensions are obtained, where the area law [6, 7] is manifested. In section IV, we provide examples in CFT2 and CFT4. In section V, we show how RT formula [8, 9] naturally emerges in this framework and discuss arbitrary shape of entangling surfaces. Section VI is the conclusion.
@@ II Entanglement entropy between two disjoint balls
Figure 3: The Euclidean pure state density matrix $\rho=|\psi\rangle\langle\psi|$ for the solid torus CFTD. The lower half solid torus corresponds to $|\psi\rangle$. The upper half annulus corresponds to $\langle\psi|$. We calculate the entanglement entropy between two disjoint co-dimensional one balls AA and $B$ (colored regions) in the bounded region $\mathcal{B}$ (blue region).
Let us consider the CFTD on a solid torus,
$\mathcal{B}=\mathbb{B}^{D-1}\times S^{1}:=\left\{(t_{\text{E}},y,x)|\left(\sqrt{t_{\text{E}}^{2}+y^{2}}-\frac{l_{2}+l_{1}}{2}\right)^{2}+\sum_{i=1}^{D-2}x_{i}^{2}<\left(\frac{l_{2}-l_{1}}{2}\right)^{2}\right\},$ (1)
where $\mathbb{B}^{D-1}$ is a (D1)(D-1)-dimensional ball and $l_{2}>l_{1}>0$. At the time slice $t_{\text{E}}=0$, $\mathcal{B}$ consists of two disjoint (D1)(D-1)-balls AA and $B$ with the same radius, as shown in Figure 3. The lower half solid torus corresponds to $|\psi\rangle$, and the upper half annulus corresponds to $\langle\psi|$. The density matrix of this pure state is $\rho=|\psi\rangle\langle\psi|$.
Using the replica trick [1], the entanglement entropy $S_{\text{disj}}(A:B)$ between $A$ and $B$ can be defined through the Rényi entropy as
$S^{(n)}(A:B)=\frac{1}{1-n}\log\mathrm{Tr}_{A}\rho_{A}^{n}=\frac{1}{1-n}\log\left[\frac{Z_{\mathcal{B}_{n}}}{Z^{n}}\right],$
$S_{\text{disj}}(A:B)=\lim_{n\to 1}S^{(n)}(A:B),$ (2)
where $\rho_{A}=\mathrm{Tr}_{B}\rho$ is the reduced density matrix. $Z$ is the partition function on the solid torus $\mathcal{B}$. $Z_{\mathcal{B}_{n}}$ is the partition function on the $n$-sheeted cover $\mathcal{B}_{n}$ made by cyclically gluing $n$ copies of $\mathcal{B}$ along $A$ (or equivalently $B$), see Figure 4 where the cut $A$ is a (D1)(D-1)-dimensional ball $\mathbb{B}^{D-1}$.
Figure 4: The cut-and-glue procedure in the replica trick to construct $\mathcal{B}_{n}$. Each solid torus is cut along the (D1)(D-1)-dimensional ball $A$ and is glued with others cyclically. Red lines represent gluing operations. Note that the resulted manifold is also a solid torus with period $2n\pi$.
Since $\mathcal{B}_{n}$ is still a solid torus with period $2n\pi$, calculating the $S_{\text{disj}}(A:B)$ in equation (2) is simplified to computing the partition function on a solid torus $\mathcal{B}=\mathbb{B}^{D-1}\times S^{1}$! A brute force approach to this calculation would involve solving the equations of motion (EoM) with either Dirichlet or Neumann boundary conditions which makes no difference since the vacuum entanglement is concerned. Instead, using the fact that a periodic system is equivalent to a thermal system, by noting the solid torus $\mathcal{B}$ is periodic, we adopt a symmetry based method: this approach is universal across all CFTs, applicable to free or strongly coupled systems.
Applying the coordinate transformation $t_{\text{E}}=r\sin\theta$, $y=r\cos\theta$ to the flat metric $\mathrm{d}s_{0}^{2}=\mathrm{d}t_{\text{E}}^{2}+\mathrm{d}y^{2}+\underset{i=1}{\overset{D-2}{\sum}}\mathrm{d}x_{i}^{2}$, we get a conformally flat metric
$\mathrm{d}s_{0}^{2}=r^{2}\left[\mathrm{d}\theta^{2}+\Big(\mathrm{d}r^{2}+\sum_{i=1}^{D-2}\mathrm{d}x_{i}^{2}\Big)/r^{2}\right].$
Removing the conformal factor $r^{2}$ by the Weyl invariance, the spacetime metric becomes
$\mathrm{d}s^{2}=\mathrm{d}\theta^{2}+\left(\mathrm{d}r^{2}+\sum_{i=1}^{D-2}\mathrm{d}x_{i}^{2}\right)/r^{2}.$ (3)
Since $\theta$ plays a role of the imaginary time with a periodicity $2\pi$, this CFT is equivalent to a thermal system with the inverse temperature $\beta=2\pi$. It is straightforward to write down the partition function of this thermal system
$Z=\mathrm{Tr}\,e^{-2\pi H},$ (4)
where the Hamiltonian operator $H$ is conjugate to the imaginary time $\theta$, defined by
@@
where $T$ is the stress-energy tensor. By noting the periodicity of the $n$-sheeted cover $\mathcal{B}_{n}$ is $2n\pi$, we immediately get the partition function on $\mathcal{B}_{n}$,
@@
Then, the Rényi entropy in equation (2) becomes
$S^{(n)}(A:B)=\frac{1}{1-n}\log\left[\frac{e^{-2\pi\langle H\rangle/n}}{e^{-2n\pi\langle H\rangle}}\right]=-2\pi\left(1+\frac{1}{n}\right)\langle H\rangle.$ (8)
The entropy $S_{\text{disj}}(A:B)$ of entanglement between two disjoint balls $A,B$ therefore reads
$S_{\text{disj}}(A:B)=S^{(1)}(A:B)=-4\pi\langle H\rangle.$ (9)
@@
The entanglement entropy $S_{\text{disj}}(A:B)$ simply depends on the ground state energy $\langle H\rangle$ localized in a spatial ball $\mathbb{B}^{D-1}$:
@@
where $\mathcal{E}_{\text{vac}}$ is the Casimir energy density that is usually a negative constant in our discussions. The integration is performed with the metric of equation (3). Consequently, the entanglement entropy $S_{\text{disj}}(A:B)$ is given by
$S_{\text{disj}}(A:B)$ $=$ $-4\pi\mathcal{E}_{\text{vac}}\mathrm{Vol}(\mathbb{B}^{D-1}),$ (11)
@@
with the standard hypergeometric function $_{2}F_{1}$. The parameters $l_{1}$ and $l_{2}$ are the inner and outer radii of the solid torus.
Figure 5: Entangling regions (colored regions) are bounded by (D2D-2)-spheres (black circles), respectively. Left panel (the cavity configuration): A (D1D-1)-ball living in a spheric cavity is entangled with the region outside the cavity. Right panel (the juxtaposed configuration): Two disjoint (D1D-1)-balls with different radii are entangled.
Generally, there exist two different configurations of our interest, namely, the cavity configuration and the juxtaposed configuration illustrated in Figure 5. Each configuration is uniquely fixed by boundaries of $A,B$, i.e., two disjoint (D2D-2)-spheres. Every (D1D-1)-ball is bounded by a (D2D-2)-sphere $\Sigma(\vec{x},r)$ located at $\vec{x}\in\mathbb{R}^{D-1}$ with the radius $r$, respectively. An elegant substitute for conformality is the inversive product $\varrho$ of two spheres $\Sigma(\vec{x},r)$ and $\Sigma(\vec{x}^{\prime},r^{\prime})$ [10]
@@
In the case that two disjoint (D1D-1)-balls $A,B$ share the same radius, we have
@@
The inversive product $\varrho$ is a conformal invariant, as it remains invariant under global conformal transformations. If two configurations have different inversive products, they cannot be conformally mapped onto each other. The set of all configurations falls into classes of conformally equivalent regions, with each class characterized by the inversive product $\varrho$ of that class. Consequently, any two configurations sharing the same $\varrho$ must be conformally equivalent. To compute the entanglement entropy $S_{\text{disj}}(A:B)$ for any two disjoint spheres in general configurations, it suffices to evaluate the inversive product $\varrho$ of the spheres. Thus, we conclude that the entanglement entropy between any two disjoint spheres is given by
$S_{\text{disj}}(A:B)=-4\mathcal{E}_{\text{vac}}\pi^{\frac{D}{2}}\frac{\Gamma\left(\frac{D}{2}\right)}{\Gamma\left(D\right)}\left(\frac{2\sqrt{2}}{\sqrt{\varrho+1}-\sqrt{2}}\right)^{D-1}\,_{2}F_{1}\left(D-1,\frac{D}{2};D;\frac{2\sqrt{2}}{\sqrt{2}-\sqrt{\varrho+1}}\right).$ (14)
The time dependent covariant scenario is obtained by simply replacing $\vec{x}$ with $x^{\mu}$.
@@ III Emergent area law in various dimensions
Figure 6: Left panel: The time zero slice of $\mathcal{B}$ in $D$ dimensions, where $A$ and $B$ are two disjoint (D1D-1)-balls with the same radius (colored regions) living at $t_{\text{E}}=0$. Right panel: The (D1D-1)-ball AA with the radius $R-\epsilon$ is in a spheric cavity with the radius $R+2\epsilon$, and the region $B$ is outside the cavity. Here, $\epsilon$ plays a role of the UV cutoff.
We first consider the entanglement entropy $S_{\mathrm{disj}}(A:B)$ for two disjoint (D1D-1)-balls $A,B$ with the same radius in the juxtaposed configuration, as shown in Figure 6. It would be convenient to write down $S_{\text{disj}}(A:B)$ in various dimensions:
@@
Then, we return to the general $D$-dimensional cavity configuration depicted in the right panel of Figure 6, where $A$ is almost adjacent to $B$, more precisely
@@
Plugging $\varrho$ into the general expression (14), we obtain $S_{\text{adj}}(A:B)$ in various dimensions:
@@
$D=6:$ $\frac{\pi^{3}\mathcal{E}_{6}}{3}\left(\frac{8R^{4}}{\epsilon^{4}}-\frac{24R^{2}}{\epsilon^{2}}+9+12\log\frac{2R}{\epsilon}\right)+\mathcal{O}\left(\epsilon\right),$ (18)
which perfectly agrees with holographic predictions given in [8]. In $D=2$, the result is the well-known $\log$-law for entanglement entropy. In $D\geq 3$, we can immediately observe the area law of entanglement entropy, which arises as a special limit of our general consideration. We emphasize that in even dimensions, the finite term is physically irrelevant since they can be absorbed into the UV cutoff $\epsilon$. But in odd dimensions, finite terms of entanglement entropies become scheme-independent and play a crucial role in the F-theorem [11, 12, 13]. This is manifested by the exactness of our expressions for entanglement entropy in odd dimensions.